# A Survey on Large Language Models for Critical Societal Domains: Finance, Healthcare, and Law

### v1

@@ 4.2 LLMs for Medicine and Healthcare
Close-sourced Medical LLMs. Close-sourced LLM pretraining for general proposes, such as ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023), have shown strong medical capacity across both medical benchmarks and real-world applications. Liévin et al. (2023) utilized GPT-3.5 with different prompting strategies, including Chain-of-thought, few-shot, and retrieval augmentation, for three medical reasoning benchmarks, to show the model’s strong medical reasoning ability in the absence of specific fine-tuning. The evaluation of LLMs, such as ChatGPT, on medical exams, including US Medical Exams (Kung et al., 2023) and Otolaryngology–Head and Neck Surgery Certification Examinations (Long et al., 2023), indicates that they achieve scores close to or at the passing threshold. This suggests the potential of LLMs to support real-world medical usages such as medical education and clinical decision-making. Agrawal et al. (2022) views LLMs such as GPT-3 as clinical information extractors and shows the potential in different information extraction tasks. The MedPaLM models (Singhal et al., 2022; 2023), are a series of medical domain-specific LLMs, adapted from PaLM models(chung2022scaling; Anil et al., 2023; Chowdhery et al., 2022), which have shown performance in answering medical questions on par with that of medical professionals (Singhal et al., 2022; 2023). GPT-4 (OpenAI, 2023) demonstrates strong medical capacities without specialized training strategies in the medical domain or engineering for solving clinical tasks (Nori et al., 2023a; b). When the scope is narrowed down to sub-domain domains, the performance of LLMs is variable. GPT-4 outperforms or performs on par with the current SOTA radiology models (Liu et al., 2023b) across radiology tasks, and it matches human performance on gastroenterology board exam self-assessments (Ali et al., 2023). Peng et al. (2023) examines ChatGPT and GPT-4 performance on Physical Medicine and Rehabilitation, and demonstrates their potential capabilities in the submedical field. However, in dementia, LLMs fail to surpass the performance of traditional AI tools (Wang et al., 2023f). Besides directly utilizing LLMs for different medical tasks, they also can be used as a knowledge base for "retrieving" informative context to auxiliary downstream tasks (Yu et al., 2023d). For example, Zhang et al. (2023g) employs ChatGPT as a medical knowledge base to generate medical knowledge for supporting downstream medical decision-making. Kwon et al. (2024) generates clinical rationales for patient descriptions and utilizes the rationale as an additional training signal to fine-tune student models in both unimodal and multimodal settings to improve diagnosis prediction performance.
@@
Open-Sourced Medical LLMs. Due to privacy concerns (Zhang et al., 2023g) and high costs, several open-source medical LLMs (Xu et al., 2023; Han et al., 2023; Li et al., 2023i; Wu et al., 2023b; Zhang et al., 2023i) have be built by tuning open-source base model, such as LLaMA (Touvron et al., 2023a; b), on medical corpus. These works mainly employ two different strategies: 1) continue pretraining followed by instruction-finetuning (Wu et al., 2023b; Xie et al., 2024a) and 2) direct instruction-finetuning (Xu et al., 2023; Han et al., 2023; Li et al., 2023i; Zhang et al., 2023i; Tran et al., 2023), as shown in 3. Specifically, the first approach involves continuously pretraining language models on biomedical corpora, including medical academic papers and textbooks, and then further fine-tuning the models with various medical instructional datasets to align with human intent for medical applications. The second approach directly conducts instruction fine-tuning on the base models to elicits the medical capabilities of base models directly. In the pretraining and fine-tuning schema, PMC-LLAMA (Wu et al., 2023b) employs a two-step training process that first extends LLaMA’s training with millions of medical textbooks and papers, and then instructionally tunes the model on a dataset of 202 million tokens. Me LLaMA (Xie et al., 2024a) proposes a domain-specific base model by continuing to pretrain LLaMA-2 models with 13B and 70B parameters on a 129 billion token medical dataset, and then creates corresponding chat models by conducting instruction fine-tuning on 219k instances. On the other hand, ChatDoctor (Li et al., 2023i) collects 100k real online patient-doctor conversations and directly fine-tunes LLaMA on the dialogues dataset. MedAlpaca (Han et al., 2023) increases the instructional dataset to 230k including question-answer pairs and dialogues and conducts the fine-tuning procedure. Baize-Healthcare (Xu et al., 2023) employs about 100k dialogues from Quora and MedQuAD for instructional tuning by LoRA (Hu et al., 2022). AlpaCare (Zhang et al., 2023i) proposes a 52k diverse machine-generated medical instructional dataset, by distilling medical knowledge from robust closed-sourced LLMs (Li et al., 2022b). Then they fine-tune open-sourced LLMs on the dataset to show the importance of training data diversity on the model’s ability to follow medical instructions while maintaining generalizability. BioInstruct (Tran et al., 2023) utilize GPT-4 to generate a 25k instruction dataset covering the topic in question-answering, information extraction and text generation to instruction tune LLaMA models (Touvron et al., 2023a; b) with LoRA (Hu et al., 2022). Their experimental results show consistent improvement in different medical NLP tasks compared to LLMs without instruction fine-tuning.
Model Name Training Schema Training Corpus Parameters
PMC-LLAMA (Liu et al., 2021) Pre-training & instruction fine-tuning 4.8M biomedical academic papers,30K textbook & 202M tokens instructional dataset. 7B, 13B
Me LLaMA (Xie et al., 2024a) Pre-training & instruction fine-tuning 129B tokens medical data & 214k medical instruction samples 13B, 70B
ChatDoctor (Li et al., 2023i) Instruction fine-tuning 100k real online patient-doctor conversations 7B
MedAlpaca (Han et al., 2023) Instruction fine-tuning; LoRA 230k data covering dialogues, online Q&A and benchmarks 7B,13B, 30B
Baize-Healthcare (Xu et al., 2023) LoRA General & 100k dialogues from Quora and MedQuAD 7B
@@
AlpaCare (Zhang et al., 2023i) Instruction fine-tuning 52k GPT-4, ChatGPT generated medical dataset 7B, 13B
BioInstruct (Tran et al., 2023) LoRA 25k GPT-4 generated medical dataset in Q&A, IE and text generation 7B, 13B
@@
Multimodal Medical LLMs. Though LLMs have the potential to process and assist clinical NLP tasks, multimodal data (e.g., X-ray Radiology, CT, MRI, ultrasound) plays a vital role in medical and healthcare applications. When making diagnosis and medical suggestions, it is important for the model to have access to and being capable of understanding clinical modalities beyond text. Hence, there is a strong need to build multi-modal LLMs (Zhu et al., 2022; Liu et al., 2024b; Yan et al., 2023a; 2024a) that are capable of connecting language with other modalities. A representative open-domain model is GPT-4V (Achiam et al., 2023), where researchers have explored its potential for understanding X-rays (Yang et al., 2023b; Wu et al., 2023a). LLaVA-Med (Li et al., 2024a) leverages a figure-caption dataset extracted from PubMed Central, and uses GPT-4 to self-instruct open-ended instruction-following data from the captions to train a visual AI assistant. Gao et al. (2023a) uses a similar recipe to train a multimodal LLM for ophthalmology.
@@ 4.3 Abnormality and Ambiguity Detection
Table 4: Evaluation (accuracy) over two categories of PLMs on abnormality identification and ambiguity identification tasks (sentence-level NLU). Bold: the highest performance. Underlined: the lowest. Results are sourced from (He et al., 2023c).
@@ 4.5 Medical Free-form Instruction Evaluation
Free-form instruction evaluations assess the practical medical value of language models from a user-centric perspective. This task involves inputting a medical query in a free-text format into the model, which then generates a corresponding response. For instance, if a user inputs, ‘Discuss the four major types of leukocytes and their roles in the human immune system in bullet point format,’ the model will produce an informed answer based on its internal medical knowledge. This task serves to measure both the medical knowledge capacity and the instruction-following capability of the model. iCliniq (Li et al., 2023i; Chen et al., 2024) contains 10k real online conversations between patients and doctors to evaluate models’ medical instruction-following ability in the dialog scenario. MedInstruct-test (Zhang et al., 2023i) contains 217 clinical craft free-form instructions to evaluate the medical capacity and instruction-following ability of models across different medical settings such as treatment recommendation, medical education, disease classification, etc. Evaluating the instruction-following capacity of LLMs is complex due to the wide range of valid responses to a single instruction and the difficulty of replicating human evaluations. Recently, automated evaluation (Zheng et al., 2023; Dubois et al., 2023; Zhang et al., 2023h; Lu et al., 2022) has offered greater scalability and explainability compared to human studies. A strong LLM is used as a judge to compare the outputs of the evaluated model with reference answers and then calculate the winning rate of the evaluated model against the reference answer is used as the evaluation metric. Table 6 shows the current open-source medical LLM performance on medical free-form instruction evaluation with GPT-3.5-turbo acts as a judge and Text-davinci-003, GPT-3.5-turbo, GPT-4, and Claude-2 as reference models, respectively.
@@
Table 6: Performance comparison of medical LLMs on medical free-form instruction evaluation. GPT-3.5-turbo acts as a judge for pairwise auto-evaluation. Each instruction-tuned model is compared with 4 distinct reference models: Text-davinci-003, GPT-3.5-turbo, GPT-4, and Claude-2. ‘AVG’ denotes the average performance score across all referenced models in each test set. The table is sourced from (Zhang et al., 2023i).
@@ 4.7 Future Prospects
  • Data Diversity and Quality: Although machine-generated datasets accelerate data generation for LLM training, their diversity still lags behind that of real-world collected datasets, which highly impacts model performance (Chiang et al., 2023). Expanding the training datasets to include a broader range of real-world medical texts, such as clinical trial reports, medical journals, patient records, and health forums, can improve the model’s understanding of diverse medical contexts and terminologies. Additionally, ensuring the quality and reliability of the training data is crucial for maintaining the accuracy and trustworthiness of the model’s outputs.
@@
  • Retrieval-Augmented Generation: Integrating retrieval-augmented generation (RAG) techniques can enhance the model’s ability to access and incorporate relevant medical knowledge from extensive sources such as large medical knowledge bases, private hospital records, and databases. This approach can improve the model’s responses by providing more accurate and contextually appropriate information, particularly in complex medical scenarios in inference time.
@@
  • Addressing Privacy Concerns: Medical data usage has strong restrictions compared to the general domain and other small domains. Developing methodologies address privacy issues in utilizing LLM APIs and building local LLMs are both important. This can include implementing secure data transmission protocols, ensuring data anonymization, and adopting privacy-preserving techniques such as differential privacy.
@@
Learning in a Data Sparsity Setting. A critical challenge in medical domain for training large-scale models is the restriction of data usage. Data sparsity is a persistent issue due to privacy and confidentiality concerns, the cost of data Acquisition and annotation, as well as ethical considerations. For many practical tasks, e.g., medical report generation, clinical chatbots, medical image classification, the data sparsity issue will be a remaining challenge. Additionally, as mentioned earlier in the performance comparison sections for different applications, task-specific models that trained with in-distribution data and specific architectural design can still outperform foundation models. Based on the empirical finding of training general domain large language models (Brown et al., 2020; Kaplan et al., 2020; OpenAI, 2023), scaling up data is of vital importance for model performance. We discuss some of the potential future directions to address this issue:
@@
  • Transfer Learning and Domain Adaptation: For medical LLMs, it is worth exploring how scaling up general domain, publicly-available data can help with in-domain medical tasks. We can explore data selection strategies in pretraining stage to improve the transfer learning performance from general-domain models to medical-specific tasks.
@@
  • Synthetic Data Generation: To mitigate the challenges posed by data scarcity, another approach is the generation of synthetic medical data. Leveraging advanced LLMs could enable the creation of diverse synthetic datasets to augment the learning process.
@@
  • Few-shot and Zero-shot Learning: Few-shot learning and in-context learning Wei et al. (2022b) methods should be explored more deeply, which has the potential to let medical LLMs adapting to new tasks or domains with minimal training data.
@@
  • Privacy-preserving Techniques: Techniques such as differential privacy and federated learning (Rieke et al., 2020) could allow the utilization of patient data for training purposes while ensuring that individual privacy is maintained.
@@
Active Learning: Implementing active learning strategies where the model identifies the most informative data points for labeling can optimize the training process. This method ensures efficient use of scarce data resources and improves learning outcomes in highly specialized medical contexts.
@@
Evaluating Real-world Medical Application Capacity. Although various benchmarks have been proposed in the medical domain, most of them focus on evaluating models from the perspective of medical knowledge (Pal et al., 2022; Jin et al., 2020; 2019), rather than from a user-oriented perspective. To bridge this gap, free-form instruction evaluation datasets utilize medical dialogues and machine-generated text for medical questions. However, these test sets are still limited in quantity and task diversity. For example, Med-Instruct proposes an evaluation dataset with diverse topics, but it is limited to only two hundred tests. On the other hand, iCliniq contains 10,000 instances, but its scope is restricted to doctor-patient conversations. Therefore, there is a need for a large-scale, diverse, and expert-verified dataset for evaluating the medical capacity of LLMs in real-world medical user applications. To evaluate the response quality of LLM in medical free-form instruction evaluation, Zhang et al. (2023i) utilizes the LLM APIs as judges (Zheng et al., 2023; Dubois et al., 2023; Zhang et al., 2023h). However, calling LLM APIs for evaluation is costly. Therefore, training a smaller LLM with strong medical capacity for response evaluation and comparison could be a more efficient alternative. Future work could focus on techniques such as knowledge distillation or model pruning to create such a medical specialized evaluator, potentially leading to faster and more cost-effective evaluation processes for medical LLM applications.

### v2

@@ 4.2 LLMs for Medicine and Healthcare
Close-sourced Medical LLMs. Close-sourced LLM pretraining for general proposes, such as ChatGPT (OpenAI, 2022) and GPT-4 (Achiam et al., 2023), have shown strong medical capacity across both medical benchmarks and real-world applications. Liévin et al. (2023) utilized GPT-3.5 with different prompting strategies, including CoT, few-shot, and retrieval augmentation, for three medical reasoning benchmarks, to show the model’s strong medical reasoning ability in the absence of specific fine-tuning. The evaluation of LLMs, such as ChatGPT, on medical exams, including US Medical Exams (Kung et al., 2023) and Otolaryngology–Head and Neck Surgery Certification Examinations (Long et al., 2023), indicates that they achieve scores close to or at the passing threshold. This suggests the potential of LLMs to support real-world medical usages such as medical education and clinical decision-making. Agrawal et al. (2022) views LLMs such as GPT-3 as clinical information extractors and shows the potential in different information extraction tasks. The MedPaLM models (Singhal et al., 2022; 2023), are a series of medical domain-specific LLMs, adapted from PaLM models(chung2022scaling; Anil et al., 2023; Chowdhery et al., 2022), which have shown performance in answering medical questions on par with that of medical professionals (Singhal et al., 2022; 2023). GPT-4 (Achiam et al., 2023) demonstrates strong medical capacities without specialized training strategies in the medical domain or engineering for solving clinical tasks (Nori et al., 2023a; b). When the scope is narrowed down to sub-domain domains, the performance of LLMs is variable. GPT-4 outperforms or performs on par with the current SOTA radiology models (Liu et al., 2023b) across radiology tasks, and it matches human performance on gastroenterology board exam self-assessments (Ali et al., 2023). Peng et al. (2023) examines ChatGPT and GPT-4 performance on Physical Medicine and Rehabilitation, and demonstrates their potential capabilities in the submedical field. However, in dementia, LLMs fail to surpass the performance of traditional AI tools (Wang et al., 2023f). Besides directly utilizing LLMs for different medical tasks, they also can be used as a knowledge base for "retrieving" informative context to auxiliary downstream tasks (Yu et al., 2023d). For example, Zhang et al. (2023h) employs ChatGPT as a medical knowledge base to generate medical knowledge for supporting downstream medical decision-making. Kwon et al. (2024) generates clinical rationales for patient descriptions and utilizes the rationale as an additional training signal to fine-tune student models in both unimodal and multimodal settings to improve diagnosis prediction performance.
@@
Open-Sourced Medical LLMs. Due to privacy concerns (Zhang et al., 2023h) and high costs, several open-source medical LLMs (Xu et al., 2023; Han et al., 2023; Li et al., 2023i; Wu et al., 2023b; Zhang et al., 2023j) have be built by tuning open-source base model, such as LLaMA (Touvron et al., 2023a; b), on medical corpus. These works mainly employ two different strategies: 1) continue pretraining followed by instruction-finetuning (Wu et al., 2023b; Xie et al., 2024a) and 2) direct instruction-finetuning (Xu et al., 2023; Han et al., 2023; Li et al., 2023i; Zhang et al., 2023j; Tran et al., 2023), as shown in 3. Specifically, the first approach involves continuously pretraining language models on biomedical corpora, including medical academic papers and textbooks, and then further fine-tuning the models with various medical instructional datasets to align with human intent for medical applications. The second approach directly conducts instruction fine-tuning on the base models to elicits the medical capabilities of base models directly. In the pretraining and fine-tuning schema, PMC-LLAMA (Wu et al., 2023b) employs a two-step training process that first extends LLaMA’s training with millions of medical textbooks and papers, and then instructionally tunes the model on a dataset of 202 million tokens. Me LLaMA (Xie et al., 2024a) proposes a domain-specific base model by continuing to pretrain LLaMA-2 models with 13B and 70B parameters on a 129 billion token medical dataset, and then creates corresponding chat models by conducting instruction fine-tuning on 219k instances. On the other hand, ChatDoctor (Li et al., 2023i) collects 100k real online patient-doctor conversations and directly fine-tunes LLaMA on the dialogues dataset. MedAlpaca (Han et al., 2023) increases the instructional dataset to 230k including question-answer pairs and dialogues and conducts the fine-tuning procedure. Baize-Healthcare (Xu et al., 2023) employs about 100k dialogues from Quora and MedQuAD for instructional tuning by LoRA (Hu et al., 2022). AlpaCare (Zhang et al., 2023j) proposes a 52k diverse machine-generated medical instructional dataset, by distilling medical knowledge from robust closed-sourced LLMs (Li et al., 2022b). Then they fine-tune open-sourced LLMs on the dataset to show the importance of training data diversity on the model’s ability to follow medical instructions while maintaining generalizability. BioInstruct (Tran et al., 2023) utilize GPT-4 to generate a 25k instruction dataset covering the topic in question-answering, information extraction and text generation to instruction tune LLaMA models (Touvron et al., 2023a; b) with LoRA (Hu et al., 2022). Their experimental results show consistent improvement in different medical NLP tasks compared to LLMs without instruction fine-tuning.
Model Name Model Architecture Training Corpus Size
MedAlpaca (Han et al., 2023) LLaMA QA rephrased from wikidoc & Flashcards, Stackexchange medical Q&A 7B,13B
ChatDoctor (Li et al., 2023i) LLaMA real patient-doctor conversations 7B
Doctor GLM (Xiong et al., 2023) GLM medical Dialogues 6.2B
Baize-Healthcare (Xu et al., 2023) LLaMA dialogues from Quora and MedQuAD 7B
@@
AlpaCare (Zhang et al., 2023j) LLaMA Alpaca data, ChatGPT & GPT-4 generated instruction data 7B, 13B
BioInstruct (Tran et al., 2023) LLaMA & LLaMA-2 GPT-4 generated instruction data 7B, 13B
Clinical Camel (Toma et al., 2023) LLaMA-2 Dialogues, articles, Medical QA 13B, 70B
BioMedGPT (Luo et al., 2023) LLaMA-2 BioMed Articles, PubChemQA, UniProtQA 7B, 10B
Meditron (Chen et al., 2023d) LLaMA-2 PubMed articles, abstracts, medical guidelines 70B
PMC-LLAMA (Wu et al., 2024) LLaMA biomedical academic papers , textbook, medical QA, rationales, dialogues 7B, 13B
Me LLaMA (Xie et al., 2024a) LLaMA medical data, medical instruction samples 13B, 70B
BioMistral (Labrak et al., 2024) Mistral PubMed Central 7B
LLaVA-Med (Li et al., 2024a) LLaVA figure-caption data from PubMed Central GPT-4 generated instruction data 7B, 13B
@@
Multimodal Medical LLMs. Though LLMs have the potential to process and assist clinical NLP tasks, multimodal data (e.g., X-ray Radiology, CT, MRI, ultrasound) plays a vital role in medical and healthcare applications. When making diagnosis and medical suggestions, it is important for the model to have access to and being capable of understanding clinical modalities beyond text. Hence, there is a strong need to build multi-modal LLMs (Zhu et al., 2022; Liu et al., 2024b; Yan et al., 2023a; 2024a) that are capable of connecting language with other modalities. A representative open-domain model is GPT-4V (Achiam et al., 2023), where researchers have explored its potential for understanding X-rays (Yang et al., 2023b; Wu et al., 2023a). LLaVA-Med (Li et al., 2024a) leverages a figure-caption dataset extracted from PubMed Central, and uses GPT-4 to self-instruct open-ended instruction-following data from the captions to train a visual AI assistant. Gao et al. (2023a) uses a similar recipe to train a multimodal LLM for ophthalmology. Zhang et al. (2023d) trained their model with image-only, text-only, and multi-modal tasks like image captioning and VQA for radiology tasks.
@@ 4.3 Abnormality and Ambiguity Detection
Table 4: Evaluation (accuracy) over two categories of PLMs on abnormality identification and ambiguity identification tasks (sentence-level NLU). Bold: the highest performance. Underlined: the lowest. Results are sourced from (He et al., 2023c). Few-shot results used 5-shots.
@@ 4.5 Medical Free-form Instruction Evaluation
Free-form instruction evaluations assess the practical medical value of language models from a user-centric perspective. This task involves inputting a medical query in a free-text format into the model, which then generates a corresponding response. For instance, if a user inputs, ‘Discuss the four major types of leukocytes and their roles in the human immune system in bullet point format,’ the model will produce an informed answer based on its internal medical knowledge. This task serves to measure both the medical knowledge capacity and the instruction-following capability of the model. iCliniq (Li et al., 2023i; Chen et al., 2024) contains 10k real online conversations between patients and doctors to evaluate models’ medical instruction-following ability in the dialog scenario. MedInstruct-test (Zhang et al., 2023j) contains 217 clinical craft free-form instructions to evaluate the medical capacity and instruction-following ability of models across different medical settings such as treatment recommendation, medical education, disease classification, etc. Evaluating the instruction-following capacity of LLMs is complex due to the wide range of valid responses to a single instruction and the difficulty of replicating human evaluations. Recently, automated evaluation (Zheng et al., 2023; Dubois et al., 2023; Zhang et al., 2023i; Lu et al., 2022) has offered greater scalability and explainability compared to human studies. A strong LLM is used as a judge to compare the outputs of the evaluated model with reference answers and then calculate the winning rate of the evaluated model against the reference answer is used as the evaluation metric. Table 6 shows the current open-source medical LLM performance on medical free-form instruction evaluation with GPT-3.5-turbo acts as a judge and Text-davinci-003, GPT-3.5-turbo, GPT-4, and Claude-2 as reference models, respectively.
@@
Table 6: Performance comparison of medical LLMs on medical free-form instruction evaluation. GPT-3.5-turbo acts as a judge for pairwise auto-evaluation. Each instruction-tuned model is compared with 4 distinct reference models: Text-davinci-003, GPT-3.5-turbo, GPT-4, and Claude-2. ‘AVG’ denotes the average performance score across all referenced models in each test set. The table is sourced from (Zhang et al., 2023j).
@@ 4.7 Future Prospects
  • Data Diversity and Quality. Although machine-generated datasets accelerate data generation for LLM training, their diversity still lags behind that of real-world collected datasets, which highly impacts model performance (Chiang et al., 2023). Expanding the training datasets to include a broader range of real-world medical texts, such as clinical trial reports, medical journals, patient records, and health forums, can improve the model’s understanding of diverse medical contexts and terminologies. Additionally, ensuring the quality and reliability of the training data is crucial for maintaining the accuracy and trustworthiness of the model’s outputs.
@@
  • RAG. Integrating RAG techniques can enhance the model’s ability to access and incorporate relevant medical knowledge from extensive sources such as large medical knowledge bases, private hospital records, and databases. This approach can improve the model’s responses by providing more accurate and contextually appropriate information, particularly in complex medical scenarios in inference time.
@@
  • Addressing Privacy Concerns. Medical data usage has strong restrictions compared to the general domain and other small domains. Developing methodologies address privacy issues in utilizing LLM APIs and building local LLMs are both important. This can include implementing secure data transmission protocols, ensuring data anonymization, and adopting privacy-preserving techniques such as differential privacy.
@@
Learning in a Data Sparsity Setting. A critical challenge in medical domain for training large-scale models is the restriction of data usage. Data sparsity is a persistent issue due to privacy and confidentiality concerns, the cost of data Acquisition and annotation, as well as ethical considerations. For many practical tasks, e.g., medical report generation, clinical chatbots, medical image classification, the data sparsity issue will be a remaining challenge. Additionally, as mentioned earlier in the performance comparison sections for different applications, task-specific models that trained with in-distribution data and specific architectural design can still outperform foundation models. Based on the empirical finding of training general domain LLMs (Brown et al., 2020; Kaplan et al., 2020; Achiam et al., 2023), scaling up data is of vital importance for model performance. We discuss some of the potential future directions to address this issue:
@@
  • Transfer Learning and Domain Adaptation. For medical LLMs, it is worth exploring how scaling up general domain, publicly-available data can help with in-domain medical tasks. We can explore data selection strategies in pretraining stage to improve the transfer learning performance from general-domain models to medical-specific tasks.
@@
  • Synthetic Data Generation. To mitigate the challenges posed by data scarcity, another approach is the generation of synthetic medical data. Leveraging advanced LLMs could enable the creation of diverse synthetic datasets to augment the learning process.
@@
  • Few-shot and Zero-shot Learning. Few-shot learning and in-context learning methods should be explored more deeply, which has the potential to let medical LLMs adapting to new tasks or domains with minimal training data.
@@
  • Privacy-preserving Techniques. Techniques such as differential privacy and federated learning (Rieke et al., 2020) could allow the utilization of patient data for training purposes while ensuring that individual privacy is maintained.
@@
Active Learning. Implementing active learning strategies where the model identifies the most informative data points for labeling can optimize the training process. This method ensures efficient use of scarce data resources and improves learning outcomes in highly specialized medical contexts.
@@
Evaluating Real-world Medical Application Capacity. Although various benchmarks have been proposed in the medical domain, most of them focus on evaluating models from the perspective of medical knowledge (Pal et al., 2022; Jin et al., 2020; 2019), rather than from a user-oriented perspective. To bridge this gap, free-form instruction evaluation datasets utilize medical dialogues and machine-generated text for medical questions. However, these test sets are still limited in quantity and task diversity. For example, Med-Instruct proposes an evaluation dataset with diverse topics, but it is limited to only two hundred tests. On the other hand, iCliniq contains 10,000 instances, but its scope is restricted to doctor-patient conversations. Therefore, there is a need for a large-scale, diverse, and expert-verified dataset for evaluating the medical capacity of LLMs in real-world medical user applications. To evaluate the response quality of LLM in medical free-form instruction evaluation, Zhang et al. (2023j) utilizes the LLM APIs as judges (Zheng et al., 2023; Dubois et al., 2023; Zhang et al., 2023i). However, calling LLM APIs for evaluation is costly. Therefore, training a smaller LLM with strong medical capacity for response evaluation and comparison could be a more efficient alternative. Future work could focus on techniques such as knowledge distillation or model pruning to create such a medical specialized evaluator, potentially leading to faster and more cost-effective evaluation processes for medical LLM applications.